OpenStudy (anonymous):

Choose the correct sum of the polynomials (4x3 − 2x − 9) + (2x3 + 5x + 3). 6x3 − 3x − 6 2x3 − 7x − 12 6x3 + 3x − 6 2x3 + 3x − 3

OpenStudy (anonymous):

6x3+3x-6

OpenStudy (anonymous):

collect simmilar terms togethr (4x3 − 2x − 9) + (2x3 + 5x + 3). =4x3 − 2x − 9 + 2x3 + 5x + 3. =4x3+2x3-2x+5x-9+3 =6x3+3x-6

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw